OpenStudy (anonymous):

Celeste made 6 cone-shaped ice blocks that each have radius of 2.8 in. and a height of 7 in. She used a cylindrical container with a radius of 9 in. to fill all of the cones. What is the height of the cylindrical container? Use 3.14 to approximate pi and round your answer to two decimal places. Show your work. Will give medal!!!!

OpenStudy (anonymous):

@bohotness

OpenStudy (bohotness):

?

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

Someone help?! @Data_LG2

OpenStudy (anonymous):

first you have to know the formula of a cone? do you know the formula?

OpenStudy (anonymous):

Yep! V=1/3pi*r^2h

OpenStudy (anonymous):

okay great :D so now you have to calculate the volume of one CONE using that formula

OpenStudy (anonymous):

i'll tell you the next step after you done with that :)

OpenStudy (anonymous):

Okay so I got 57.44106(repeating)7

OpenStudy (anonymous):

why you multiply it by 7? shouldn't it be 6?

OpenStudy (anonymous):

No the six is repeating and stops the 6s

OpenStudy (anonymous):

* 7*

OpenStudy (anonymous):

" 6 cone-shaped ice blocks "

OpenStudy (anonymous):

so it means, after you find out the value of the volume of one cone, you have to multiply it by 6... this outcome will be the volume of the cylinder

OpenStudy (anonymous):

okay so that's 344.6464

OpenStudy (anonymous):

\(\sf \large \frac{1}{3} \pi r^2h= \frac{1}{3}(3.14)(2.8)^2(7)= 9.2022\) multiply it by 6 because you have 6 cones, that will be \(\sf 9.02 \times 6 = 55.21\) this will be equal to the volume of the cylinder (: do you know the formula?

OpenStudy (anonymous):

Sorta

OpenStudy (anonymous):

The answer is 6.13 isnt it

OpenStudy (anonymous):

\(\sf V= \pi r^2h\\55.21= (3.14)(9)^2h\) solve for h \(\sf \large h=\frac{55.21}{(3.14)(9^2)}\)

OpenStudy (anonymous):

oops i made a mistake somewhere

OpenStudy (anonymous):

hmm give me a sec..

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @Data_LG2 \(\sf \large \frac{1}{3} \pi r^2h= \frac{1}{3}(3.14)(2.8)^2(7)= 9.2022\) multiply it by 6 because you have 6 cones, that will be \(\sf 9.02 \times 6 = 55.21\) this will be equal to the volume of the cylinder (: do you know the formula? \(\color{blue}{\text{End of Quote}}\) this should be \(\sf \large \frac{1}{3} \pi r^2h= \frac{1}{3}(3.14)(2.8)^2= 8.2058 \times 6=49.23\) \(\sf V= \pi r^2h\\49.23= (3.14)(9)^2h\) solve for h \(\sf \large h=\frac{49.23}{(3.14)(9^2)}=0.18\) hmm this is kind of strange....

OpenStudy (anonymous):

@confluxepic sorry i tagged you, can you help me please? tso is not online

OpenStudy (confluxepic):

@TheSmartOne

OpenStudy (confluxepic):

This is confusing.

TheSmartOne (thesmartone):

Everything looks correct. Hmm, @iambatman @Directrix @whpalmer4 Can uoi please check? Thanks :)